OpenStudy (anonymous):

Six tiles lettered C, O, O, K, I, and E are lying facedown on a table. A tile is selected, the letter is recorded, the tile is replaced, and the process is repeated. What is the theoretical probability of selecting the letter 0? 1/3 1/2 1/6 2/5

OpenStudy (anonymous):

2 O's in 6 possibles is p=2/6=1/3

OpenStudy (lilai3):

1/3, because there are 6 letters in the word COOKIE, and there are 2 "o"'s in that word. the fraction is 2/6, which is reduced to 1/3. the answer is 1/3.